# A Set of Rules for Model Validation

### v1

@@ 2.5 Rule 5: Comparisons among models should consider both statistical and practical significance
Favoring the model or data pipeline with the optimal average performance is tricky: for enough resolution there is always a ’perceived’ winner, like there is always a tallest person in the room. However, there may be cases in which models or pipelines are simply even in performance, given the level of uncertainty in the data. For instance, this happens when the validation is repeated with different tests sets and the best model changes. For this reason, rather than reporting only average performance metrics (like total sum-of-squares of the error), reporting uncertainty statistics and/or statistical significance is more informative. Taking statistical significance into account at model selection, but in particular for performance evaluation, has practical benefits, because we can always choose the most convenient solution among the similarly performing models. For performance measures that can be easily broken down at observation level (like the PRESS), we can directly use its distribution along the observations to compute uncertainty or statistical significance [^3] Care should be taken to properly manage the correlation between performance estimates within and between models. An alternative way to provide uncertainty metrics is through resampling techniques like bootstrapping [22, 23]. Uncertainty in the model performance is relevant information to be reported for transparency and future comparison.
Take now the example of Figure 6, based on simulated data where $\mathbf{X}~(20\times 100)$ and $\mathbf{y}~(20\times 2)$ have a noisy connection between $\mathbf{y}$ and a subset of variables in $\mathbf{X}$. We perform double CV with 10 repetitions for different PLS-based models, with and without variable selection (including the selectivity Ratio (SR) [24], Variable Importance in Projection (VIP) [25] and sparse PLS [26]). Now, lets assume we disregard variability information and we only consider one repetition, e.g., repetition 1. Our conclusion would be that the performance (from best to worst) is VIP-PLS, SR-PLS, sPLS, and PLS. If we rather take repetition 2, we get sPLS, SR-PLS, VIP-PLS, and PLS. And if we take repetition 9: PLS, SR-PLS, VIP-PLS, sPLS. The boxplot provides a more informative comparison, by leveraging the variability in the results: PLS and sPLS provide moderately unstable performance results and SR-PLS and VIP-PLS provide both more stable and better performance. VIP-PLS attains the maximum median, but (very importantly) there is no statistically significant difference between SR-PLS and VIP-PLS (p-value = 0.26), so both models can be considered even since the difference in perceived performance can be the result of chance.
@@
However, even when statistically significant differences exist among alternative models or data pipelines, the differences may still be negligible from the practical standpoint. Furthermore, there may be additional considerations to make when selecting the ’best’ solutions: some models may require impractical data volumes, processing times and/or energy consumption for training, or make use of impractical information complex or expensive to retrieve, like detailed data labeling for supervised models. An extended practice in literature is to propose a supervised deep learning model even in applications where a reliable data labeling may be obtained only manually. This manual labeling may be impractical or even unfeasible, especially for models that require regular updating. Furthermore, an important consideration that is often forgotten is the labeling uncertainty. In most real applications, a certain level of labeling error is always expected, and pushing accuracy above this level is of limited practical value.
In the previous example, computational times for the double CV was: 1.3s for PLS, 21.8s for SR-PLS, 16.9s for VIP-PLS and 26.3s for SPLS. Depending on the situation (rates at which models need to be created and validated, expenditures in money and energy associated, etc.) the PLS option may be a more practical solution.
@@ 3 Assessing an Example of Validation Strategy: the Double-check for PLS-DA
Szymańska et al. [27] propose a very complete validation scheme for metabolomic models with Partial Least Squares Discriminant Analysis (PLS-DA). The proposal combines double CV with permutation testing. The double CV splits the data intro three datasets: model building and test data are separated in the outer loop to guarantee independence between them, and model buiding data is further split into training and validation data in the inner loop, to fit several PLS-DA models with different numbers of LVs. The inner loop then selects an optimal model variant and the generalization performance is estimated with the independent test. The exact same process is repeated many times after permuting the rows in Y or X, leaving the other block untouched, so that a null distribution comprising the spurious relationships in the data is estimated. This is repeated multiple times to assess the uncertainty in the generalization performance estimates. We extended this algorithm to sparse PLS-DA [28, 29], with the addition that both the number of LVs and the sparsity level are to be chosen in the inner loop, and that uncertainty can be used for model selection.
@@
    Rule 4: A baseline is provided based on permutation testing, to evaluate if the generalization performance of the model is due to chance or rather statistically significant. This is a very interesting complement to the double CV loop because it allows us to determine whether a given performance result is actually significant for a dataset with specific characteristics: number of rows, number of columns, correlation within X or Y and, very importantly, imbalanced nature.
@@
    Rule 5: While not implemented in the original publication of Szymańska et al. [27], the resampling approach within the CV loop can be leveraged to estimate uncertainty measures in the jackknifing/bootstrap style  [22, 31], which can be taken into account at model selection. Furthermore, the uncertainty reflected by the repetitions of the double CV can be taken into account when comparing the generalization performance of (s)PLS-DA with other model variants (in the spirit of Figure 6). It should be considered, however, that these variance estimates are not independent and that practical significance should always be taken into account [32].
@@ 5 Acknowledgements
I kindly acknowledge the comments by Michael Sorochan Armstrong, Jesús García-Sánchez and Daniel Vallejo España. This work was supported by grant no. PID2023-1523010B-IOO (MuSTARD), funded by the Agencia Estatal de Investigación in Spain, call no. MICIU/AEI/10.13039/501100011033, and by the European Regional Development Fund.

### v2

@@ 2.5 Rule 5: Comparisons among models should consider both statistical and practical significance
Favoring the model or data pipeline with the optimal average performance is tricky. If the criterion for performance evaluation has enough resolution—i.e., enough decimal places—there is always a ’perceived’ winner. However, there may be cases in which models or pipelines are simply even in performance, given the level of uncertainty in the data. For instance, this may be the case when the validation is repeated with different tests sets and the best model changes. For this reason, rather than reporting only average performance metrics, like total sum-of-squares of the error, reporting uncertainty statistics and/or statistical significance is more informative. Taking statistical significance into account at model selection, but in particular for performance evaluation, has practical benefits, because we can always choose the most convenient solution among the similarly performing models. For performance measures that are computed per observation (like the PRESS), we can directly use the distribution of these values across the observations to compute uncertainty or statistical significance [^4] Care should be taken to properly manage the correlation between performance estimates within and between models. An alternative way to provide uncertainty metrics is through resampling techniques like bootstrapping [22, 23]. Uncertainty is a critical component of model performance and is necessary for conducting rigorous comparisons across models. Consequently, model uncertainty should be reported to ensure transparency and facilitate future benchmarking.
Take now the example of Figure 6, based on simulated data where $\mathbf{X}~(20\times 100)$ and $\mathbf{y}~(20\times 2)$ have a noisy connection between $\mathbf{y}$ and a subset of variables in $\mathbf{X}$. We perform double cross-validation with ten repetitions for different PLS-based models, with and without variable selection, including the selectivity Ratio (SR) [24], Variable Importance in Projection (VIP) [25] and sparse PLS [26]. Setting aside variability for a moment, let us focus on a single run: repetition one. If we only look at that repetition, our conclusion would be that the performance (from best to worst) is VIP-PLS, SR-PLS, sPLS, and PLS. If we rather take repetition two, we get sPLS, SR-PLS, VIP-PLS, and PLS. And if we take repetition nine: PLS, SR-PLS, VIP-PLS, sPLS. The boxplot provides a more informative comparison, by leveraging the variability in the results: PLS and sPLS provide moderately unstable performance results and SR-PLS and VIP-PLS provide both more stable and better performance. VIP-PLS attains the maximum median, but (very importantly) there is no statistically significant difference between SR-PLS and VIP-PLS (p-value = 0.26), so both models can be considered even since the difference in perceived performance can be the result of chance.
@@
However, even when statistically significant differences exist among alternative models or data pipelines, the differences may still be negligible from the practical standpoint. Furthermore, there may be additional considerations to make when selecting the ’best’ solutions: some models may require impractical data volumes/sizes, processing times and/or energy consumption for training, or make use of impractical information complex or expensive to retrieve, like detailed data labeling for supervised models. An increasingly widespread practice in recent data science literature is to propose a supervised deep learning model even in applications where reliable data labeling may be obtained only manually. This manual labeling may be impractical or even unfeasible, especially for models trained from massive data that require regular updating. Furthermore, an important consideration that is often forgotten is the labeling uncertainty. In most real applications, a certain level of labeling error is always expected, and pushing accuracy above this level is of limited practical value—it is, as a matter of fact, overfitting, but mostly a waste of time and resources.
In the previous example, computational times for the double cross-validation was: 1.3s for PLS, 21.8s for SR-PLS, 16.9s for VIP-PLS and 26.3s for SPLS. Depending on the situation—rates at which models need to be created and validated, expenditures in money and energy associated, etc.—the PLS option may be a more practical solution.
@@ 3 Assessing an Example of Validation Strategy: the Double-check for PLS-DA
Szymańska et al. [27] propose a very complete validation scheme for metabolomic models with Partial Least Squares Discriminant Analysis (PLS-DA). The proposal combines double cross-validation with permutation testing. The double cross-validation splits the data intro three datasets: model building and test data are separated in the outer loop to guarantee independence between them, and model buiding data is further split into training and validation data in the inner loop, to fit several PLS-DA models with different numbers of LVs. The inner loop then selects an optimal model variant and the generalization performance is estimated with the independent test. The exact same process is repeated many times after permuting the rows in Y or X, leaving the other block untouched, so that a null distribution comprising the spurious relationships in the data is estimated. This is repeated multiple times to assess the uncertainty in the generalization performance estimates. We extended this algorithm to sparse PLS-DA [28, 29], with the addition that both the number of LVs and the sparsity level are to be chosen in the inner loop, and that uncertainty can be used for model selection.
@@
    Rule 4: A baseline is provided based on permutation testing, to evaluate if the generalization performance of the model is due to chance or rather statistically significant. This is a very interesting complement to the double cross-validation loop because it allows us to determine whether a given performance result is actually significant for a dataset with specific characteristics: number of rows, number of columns, correlation within X or Y and, very importantly, imbalanced nature.
@@
    Rule 5: While not implemented in the original publication of Szymańska et al. [27], the resampling approach within the cross-validation loop can be leveraged to estimate uncertainty measures in the jackknifing/bootstrap style  [22, 31], which can be taken into account at model selection. Furthermore, the uncertainty reflected by the repetitions of the double cross-validation can be taken into account when comparing the generalization performance of (s)PLS-DA with other model variants (in the spirit of Figure 6). It should be considered, however, that these variance estimates are not independent and that practical significance should always be taken into account [32].
@@ 5 Acknowledgements
I kindly acknowledge the comments by Michael Sorochan Armstrong, Jesús García Sánchez and Daniel Vallejo España, and of the anonymous reviewers. All of them were really helpful to improve the message of the paper. This work was supported by grant no. PID2023-1523010B-IOO (MuSTARD), funded by the Agencia Estatal de Investigación in Spain, call no. MICIU/AEI/10.13039/501100011033, and by the European Regional Development Fund.